OpenStudy (anonymous):

Has anyone here taken "Quantum Field Theory" Class??? I need help [Reward: Medal and Fan]

OpenStudy (anonymous):

What's your question

OpenStudy (anonymous):

wait a minute . . .

OpenStudy (anonymous):

btw, i'm confused how to write dho (partial differential sysmbol) using equation here??

OpenStudy (anonymous):

rho? \[\partial\]

OpenStudy (anonymous):

`\partial`

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

Scalar charged particle is described by lagrangian \[L = \frac{ 1 }{ 2 } (\partial_\mu \phi^* \partial^\mu \phi -m^2 \phi^* \phi)\] where m is particle mass and \(\phi(x)\) is complex scalar field which satisfies an equal time commutation relation \[[\phi(x),\pi(y)] = [\phi^*(x),\pi^*(y)] = i\delta^3(\bar{x}-\bar{y})\] where x(y) is canonical momentum field

OpenStudy (anonymous):

a) Show that the lagrangian above gives \[\partial_\mu \partial^\mu \phi + m^2 \phi = 0\]

OpenStudy (anonymous):

b) Show that \[\phi(x) = \int \frac{ d^3k }{ (2\pi)^3} \frac{ 1}{ 2 \omega_k } (a_ke^{-ik.x}+ b^\dagger_{k} e^{ik.x})\] where k is four-momentum and \(\omega_k\) is particle energy, is solution of \(\partial_\mu \partial^\mu \phi + m^2 \phi = 0\)

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

c) Show that both \(a_k\) and \(b^\dagger_k\) operator satisfy commutation relation : \[[a_k,a^\dagger_q] = [b_k,b^{\dagger}_q] = (2\pi)^3 2\omega_k \delta^3 (\bar{k}-\bar{q})\]

OpenStudy (anonymous):

d) Calculate the Hamiltonian for the system: \(H = \int\limits d^3x : H : \) with \(H = \sum_i \pi_i \dot{\phi_i} - L\), where L is lagrangian

OpenStudy (anonymous):

@ParthKohli, @Compassionate, @radar, @uri, @geerky42, @ganeshie8, @Abhisar

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Sure, so what have you done?

OpenStudy (anonymous):

i have finished part a)

OpenStudy (anonymous):

have idea @Pompeii00 ?

OpenStudy (anonymous):

I've never learned this stuff, but I've read about it on my own before. Have you tried to approach the first part (a)? I'm thinking it's probably just applying the Euler-Lagrange relation:\[\frac{\partial L}{\partial (\partial_\mu \phi)}=\frac{\partial L}{\partial_\mu \phi}\]

OpenStudy (anonymous):

i have finished part a),

OpenStudy (anonymous):

\[\frac{ \partial L }{ \partial \phi^*(x) } = \frac{ \partial }{ \partial \phi^*(x) }\left( \frac{ 1 }{ 2 } \partial_\mu \phi^* \partial^\mu \phi - \frac{ m^2 }{ 2 } \phi^* \phi \right)\]

OpenStudy (anonymous):

then

OpenStudy (anonymous):

\[\frac{ \partial L }{ \partial \phi^* (x) } = -\frac{ m^2 }{ 2 } \phi\]

OpenStudy (anonymous):

and \[\frac{ \partial L }{ \partial (\partial_\mu \phi^*(x)) } = \frac{ 1 }{ 2 } \partial^\mu \phi\]

OpenStudy (anonymous):

based on euler-lagrange equation \[\frac{ \partial L }{ \partial \phi^*(x) } - \partial_\mu \frac{ \partial L }{ \partial (\partial_\mu \phi^*(x)) } =0\] \[-\frac{ m^2 }{ 2 } \phi - \partial_\mu (\frac{ 1 }{ 2 } \partial^\mu \phi ) = 0\] \[\partial_\mu \partial^\mu \phi + m^2 \phi = 0\]

OpenStudy (anonymous):

how about part b, c, and d?? i really need help

OpenStudy (anonymous):

I agree. That's a nice derivation. I'm not sure about (b). I notice a complex exponential term, which reminds me of a Fourier Transform.

OpenStudy (surry99):

nice hat

OpenStudy (michele_laino):

for answer b), please you have to substitute the solution \[\phi \] into your equation: \[(\partial ^{\mu} \partial _{\nu} -m ^{2})\phi=0\] so you will get this: \[k ^{\mu}k _{\mu}-m ^{2}=0\] Please note that your solution is in the form of Fourier transform, so it is a tipical practice, substitute it into the equation which we attempt to solve in order to find some constraint

OpenStudy (michele_laino):

so the length of 4-vector k is equal to m

OpenStudy (anonymous):

ok.., i'll try

OpenStudy (michele_laino):

oops...your equation is: \[(\partial _{\mu} \partial ^{\mu} -m ^{2})=0\]

OpenStudy (anonymous):

\[(\partial_\mu \partial^\mu + m^2) \phi(x) \] \[= (\partial_\mu \partial^\mu + m^2) \left( \int\limits \frac{ d^3k }{ (2\pi)^3 } \frac{ 1 }{ 2 \omega_k } \left( a_k e^{-ik.x} + b^\dagger_k e^{ik.x} \right)\right)\] \[=\int\limits \frac{ d^3k }{ (2\pi)^3} (\partial_\mu \partial^\mu + m^2) \frac{ 1 }{ 2 \omega-k } (a_k e^{-ik.x} + b_k^\dagger e^ik.x)\] \[= \int\limits \frac{ d^3k }{ (2\pi)^3 } \frac{ 1 }{ 2 \omega_k } \left( (-k^2 + m^2) a_k e^{-ik.x} + (-k^2+ m^2) b_k^\dagger e^{ik.x} \right)\] \[a_k = a(k)\] \[b^\dagger_k = a^\dagger(k)\] then \[= \int\limits \frac{ d^3k }{ (2\pi)^3 } \frac{ 1 }{ 2 \omega_k } \left( (-k^2 + m^2) a(k) e^{-ik.x} + (-k^2+ m^2) a^\dagger (k) e^{ik.x} \right)\]

OpenStudy (michele_laino):

I used another method, namely the tensorial calculus, but the answer is the same!

OpenStudy (anonymous):

would you mind writing it here ??

OpenStudy (anonymous):

or just take a photo

OpenStudy (michele_laino):

please, wait a moment....

OpenStudy (anonymous):

ok thank you ;)

OpenStudy (michele_laino):

I'm sorry, I tried with my computer cam, but the image is bad quality!

OpenStudy (michele_laino):

never mind, I put into integral sign the two derivations successively

OpenStudy (michele_laino):

@gerryliyana sorry, I have a little question: the operators a, namely a_k and a_q do commute?

OpenStudy (anonymous):

yes...,

OpenStudy (anonymous):

\[[a_k,a_q]=0\]

OpenStudy (anonymous):

\[[a_k,a_q] = 0 = [a_k^\dagger, a_q^\dagger]\]

OpenStudy (anonymous):

\[[a_k, a_q^\dagger] = \delta^3 (k-q)\]

OpenStudy (michele_laino):

ok! Now I try to insert two solution, namely: \[\phi (x)=\int\limits \frac{ d ^{3} k}{ (2 \pi)^{3} }\frac{ 1 }{ 2 \omega _{k} }(a _{k}e ^{-ikx}+b _{k}^{*}e ^{ikx})\] and: \[\phi (x)=\int\limits \frac{ d ^{3} k}{ (2 \pi)^{3} }\frac{ 1 }{ 2 \omega _{k} }(a _{k}e ^{-iky}+b _{k}^{*}e ^{iky})\] into your commutation relation

OpenStudy (michele_laino):

oops... the second solution is \[\pi (x)\]

OpenStudy (michele_laino):

oops.. in the second solution please read q in place of k

OpenStudy (anonymous):

ok.., and then what's the next step??

OpenStudy (michele_laino):

wait a moment please, I'm trying...

OpenStudy (anonymous):

ok ..., ;)

OpenStudy (michele_laino):

after that substitution, I get, as integrand function, the function below:

OpenStudy (michele_laino):

\[([a _{k},b _{q}^{*}]e ^{-ikx+iqy}+[b _{k}^{*},a _{q}]e ^{ikx-iqy})\frac{ 1 }{ (2 \pi)^{6} }*\frac{ 1 }{ 4\omega _{k}\omega _{q} }\]

OpenStudy (michele_laino):

where [,] is the commutator operator

OpenStudy (michele_laino):

now, we have to perform two integrations, namely \[d ^{3}q,d ^{3}k\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

now, Isince we have: \[b _{k}^{*}=a _{k}^{*}\] I use your above commutation relation between a's and between a's and a*'s

OpenStudy (michele_laino):

the first term can be rewritten as below: \[(2 \pi)^{3}\delta ^{3}(k-q)e ^{-ikx+iqy}\]

OpenStudy (michele_laino):

and the second term, can be rwritten as below: \[(2 \pi)^{3}\delta ^{3}(k-q) e ^{ikx-iqy}\]

OpenStudy (michele_laino):

Sorry, I think to have made an error sign!

OpenStudy (michele_laino):

I got: \[[\phi (x),\pi(y)]=0\]

OpenStudy (anonymous):

great!

OpenStudy (michele_laino):

I try to evaluate this commutator: \[[\phi (x),\pi ^{*}(y)]\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

I got: \[[\phi (x),\pi ^{*}(y)]=0\]

OpenStudy (michele_laino):

that's very strange, I think to have made an error, but I don't found it!

OpenStudy (michele_laino):

in other words, if I apply the commutation relation, namely: \[[a _{k},a _{q}^{*}]=\delta ^{3}(k-q)\] I lways get zero for all commutator between \[\phi (x),\pi(y)\]

OpenStudy (michele_laino):

and their complex conjugates

OpenStudy (michele_laino):

@gerryliyana what do you think about that, please?

OpenStudy (anonymous):

i absolutely agree with you @Michele_Laino ...,

OpenStudy (michele_laino):

ok! Now I try to find your commutation relations, starting from your second equation of your problem

OpenStudy (anonymous):

i'm ready . . .

OpenStudy (michele_laino):

I evaluated this commutator: \[[\phi (x),\pi (y)]\] and I got these four terms: \[[a _{k},a _{q}]e ^{-ikx-iqy}\]

OpenStudy (michele_laino):

\[[a _{k},a _{q}^{*}]e ^{-ikx+iqy}\]

OpenStudy (michele_laino):

\[[a _{k}^{*},a _{q}]e ^{ikx-iqy}\]

OpenStudy (michele_laino):

and: \[[a _{k}^{*},a _{q}^{*}]e ^{ikx+iqy}\]

OpenStudy (michele_laino):

now, we have to find necessary conditions under which that commutator is equal to: \[i \delta ^{3}(x-y)\]

OpenStudy (michele_laino):

I think that those conditions are: \[[a _{k},a _{q}]=[a _{k}^{*},a _{q}^{*}]=0\] and: \[[a _{k},a _{q}^{*}]=[a _{k}^{*},a _{q}]= constant*i \delta ^{3}(k-q)\] where i is such that: \[i ^{2}=-1\]

OpenStudy (anonymous):

i found this

OpenStudy (anonymous):

look at page 9., there is the same problem with ours

OpenStudy (michele_laino):

at that page there are solutions of some exercises, better it is if we know also texts of those exercise, I think!

OpenStudy (anonymous):

have you seen equation 3.41 at that page?

OpenStudy (anonymous):

btw thank you for helping me @Michele_Laino . I really appreciate it.

OpenStudy (michele_laino):

Thank you! @gerryliyana